OpenStudy (anonymous):

what is the area of a triangle with a base of nine units and height of 22 units?

OpenStudy (anonymous):

99 units

OpenStudy (anonymous):

Area of triangle = 1/2 x base x height = 1/2 x 9 x 22 = 198/2 = 99